# Mechanisms and Robotics

### v1

@@ 3.3 ABB GoFa CRB 15000 is Cuspidal
The ABB GoFa [31] is a new line of cobots with three kinematic variants: 5 kg, 10 kg, and 12 kg. The 5 kg and 10 kg variants have intersecting axes $(1,2)$. All variants have parallel axes $(2,3)$, intersecting axes $(4,5)$, and nonconsecutive intersecting axes $(4,6)$. We found examples of non-singular changes of IK solutions for all three variants. To the best of our knowledge, we are the first to report that the GoFa is a cuspidal manipulator.
@@ 3.4 Robots with Three Parallel Axes May Be Cuspidal
Others have mentioned that robots with three parallel axes are noncuspidal [7]. Although this is true for common robots such UR cobots which have other intersecting axes and can be solved with a series of quadratics, in genera, this is false. It is easy to find a counterexample by searching through random robots and poses and testing if any MoveJ path is nonsingular. For example, the robot with kinematic parameters
@@ 4 Path Planning Using Graph Method
While the IK problem finds the joint angles for a single end effector pose, the path planning problem extends this to find the entire joint path for an end effector path. It is not sufficient to apply IK to each point in the end effector path because the joint path must be continuous and may have other constraints, and the IK problem admits multiple solutions.
@@ 4.1 Problem Formulation
Denote an end effector path $\underline{\chi}_{0T}(\lambda)$ with path index $\lambda\in[0,L]$, and denote a point along the path by $\chi_{0T}(\lambda)$. The goal of the path planning problem is to find the optimal continuous corresponding joint-space path $\underline{q}(\lambda)$. If desired, the path can be indexed to time with a trajectory $\underline{\lambda}(t)$, where $t\in[0,T]$. We can discretize the path into $N+1$ evenly spaced points so that $L=N\Delta\lambda$. Define the discrete path index $k\in\{0,1,\dots,N\}$ so that $\chi_{0T}[k]=\chi_{0T}(k\Delta\lambda)$. Similarly, we can represent the joint-space solutions using the discretized path $\underline{q}[k]$.
@@ 4.2 Graph Method
For a 6-DOF manipulator, the task-space pose $\chi_{0T}[k]=(R_{0T}[k],p_{0T}[k])$ consists of the orientation and position of the end effector task frame in the robot base frame. For each $\chi_{0T}[k]$ the inverse kinematics solutions are $q^{1}[k],\ q^{2}[k],\ \dots,\ q^{M[k]}[k]$, where $M[k]$ is the number of IK solutions for $\chi_{0T}[k]$. (A special case is when the IK solutions include a solution at an internal singularity, in which case self-motion is possible at that solution.) We construct a weighted directed acyclic graph in the following manner:
@@
  • Each IK solution vertex $V_{k,m}$ has a weighted edge to $V_{k+1,l}$ if the metric $c(q^{l}[k+1],q^{m}[k]))$ is less than some user-defined $\epsilon=\Delta\lambda\epsilon_{0}$, where $\epsilon_{0}\approx\left\|{\dot{q}_{max}}\right\|^{2}$.
  • Vertex $S$ is connected to each $V_{0,m}$, and vertex $F$ is connected to each $V_{N,m}$ with edges of weight 0 directed from $S$ and to $F$.
@@
Limits on certain joint angles mean that we need to take into account the total number of turns rather than only considering $q_{i}\in[-\pi,\pi]$. Care must be taken for cyclical paths so that repeating the path does not cause the joint to exceed the joint limit. Velocity or acceleration constraints can be used to determine if two vertices are connected with an edge (hard constraint) or they can be incorporated into the cost of each edge (soft constraint).
@@
Paths encounter singularities if $\det(J)=0$. Only nonsingular paths are returned if edges are not made between vertices with different signs for $\det(J)$.
@@
By finding all connections between nodes at $k=0$ and $k=N$ (which represent the same IK solutions), we can make a new directed graph showing connections between IK solutions. Then, we must find graph cycles starting and ending at the same IK solution.
Alternatively, rather than constructing a new graph, we can instead connect each $V_{0,m}$ to the corresponding $V_{N,m}$ and find minimum-weight cycles directly.
@@
To account for missed IK solutions, also connect vertex $q^{m}[k]$ to vertex $q^{l}[k+2]$ if the metric
@@
is less than $\epsilon$. Similarly, $S$ and $F$ should be connected to $V_{1,m}$ and $V_{N-1,m}$. If desired, nodes more than 2 timesteps away may be connected as desired similarly.
Due to the triangle inequality, paths with fewer edges and with larger $\Delta\lambda$ will have a smaller weight than paths between the same vertices with more edges and with smaller $\Delta\lambda$. To reduce the error in calculating the total path cost, edges should be added to the graph in multiple passes starting with the smallest $\Delta\lambda$, and vertices should be connected with edges with larger $\Delta\lambda$ only if they are not already connected.
Another strategy to prevent missed IK solutions when using IK-Geo is to utilize continuous approximate solutions. This allows for Cartesian paths at or even through boundary singularities. This means paths can be at the edge of the workspace or can switch between aspects. Care should be taken to ensure the task-space path deviation is not excessive when using these solutions. Continuous approximate solutions are also useful as part of IK based on 1D search to return IK solutions at the edge of the error function branch where the function becomes vertical, in which case the robot may not be singular.
@@ 4.3 Evaluation
An example of the graph method for planning a straight-line Cartesian move for the FANUC CRX-10iA/L is shown in Fig. 8. This example uses 10 equally-spaced sample points to make the graph easy to see, but in practice, the density of sample points should be higher.
Furthermore, Figs. 2, 3, and 4 were constructed using the graph method.
@@ 5 Path Optimization
It is common for robots to perform operations such as welding where there is freedom in how to fixture the workpiece with respect to the robot, either by moving the workpiece or by moving the robot. In such a scenario, a natural problem to solve is how to find the optimal pose of the workpiece based on some optimization metric.
@@ 5.2 Solution Method
A locally optimal toolpath offset $\chi$ can be found using an optimization method such as fminsearch() in MATLAB [33], which is a derivative-free local optimizer. To find a feasible starting point, randomly search over $\mathrm{SO}(3)$ and $\mathbb{R}^{3}$ based on the robot workspace until the graph-based path planner finds a feasible $q[k]$. Then, optimization is performed with the graph-based path planner running at each iteration. If desired, the global minimum can be found with high probability if the optimization is performed with multiple random starting points.
Represent the rotation $R$ by a unit quaternion with scalar part $\cos(\theta/2)$ and vector part $k\sin(\theta/2)$, where $k$ is the axis of rotation and $\theta$ is the rotation angle. During optimization, the unit-length constraint does not need to be enforced and the quaternion can be normalized before use, although it should be verified after optimization that the norm does not become too small or large.
@@
If there are no joint limits or position constraints that depend on $q_{1}$ then rotating the path about the first joint axis $h_{1}$ does not change the optimization.
@@
It remains to show that any rotation $R$ can be represented as a rotation about a vector in the $xy$ plane followed by a rotation about $z$, that is $R=R_{z}R_{xy}$. Decompose the rotation into $z$$x$$z$ Euler angles:
@@ 5.3 Evaluation
Optimization was performed for both a 3R and 6R robot using a helical toolpath. The optimization metric was the $L_{2}$ joint velocity norm approximation without any other constraints.
@@
Optimization was also performed for the CRX-10iA/L with a 6-DOF helical toolpath and two initial guesses. Results are shown in Figs. 9(b), 10(b), and 11(b). RMS joint velocity over path length decreased from 13.9328 rad/m to 10.0447 rad/m for the optimized pose near $z=0.75$ and decreased from 13.1054 rad/m to 9.9929 rad/m for the optimized pose near $z=-0.5$.
@@ 6 Conclusion
Most robot practitioners are unaware of cuspidal robots, and many of those who are aware suggest cuspidal robots should be avoided. Our suggestion is that cuspidal robots are difficult to use not because of an inherent problem with the kinematics but because there is a current lack of effective IK and path planning algorithms to control them. Indeed, cuspidal robots offer certain advantages. Wrist offsets may lead to simpler mechanical designs and better dynamical performance. Such designs even allow for the upper arm to swing past the lower arm, such as for some models in the CRX line. Having multiple IK solutions available in a single aspect can be beneficial because one IK solution may be preferred over another for obstacle or singularity avoidance, and switching between solutions does not require passing through a singularity.
Therefore, it is important not to dismiss cuspidal robots due to their path planning issues but instead to develop new path planning tools so that cuspidal robots can be used to their full potential. The work presented in this paper is one of the few first steps in improving the path planning capabilities of cuspidal robots. We have shown how to identify cuspidal robots, how to perform path planning given a task-space path, and how to perform path optimization. For the first time, we have shown the ABB GoFa and certain robots with three parallel joint axes are cuspidal. The algorithms presented underscore the importance of IK solvers which can efficiently find all IK solutions, such as IK-Geo.
@@
  • What is a good metric for the “degree of cuspidality" for a robot? It is apparent that different robots have different proportions of random poses with non-singular MoveJ changes of IK solution (I.e., the algorithm in Section 3.2 may have a different expected iteration count) and different proportions of random MoveL paths that are infeasible. Such a metric may even be useful for noncuspidal robots.
@@
This work in identification, path planning, and path optimization addresses critical theoretical challenges in cuspidal robots while improving their practical viability. Advancing these methods and addressing the open questions we raise will further unlock the full potential of cuspidal manipulators.

### v2

@@ 3.3 ABB GoFa CRB 15000 is Cuspidal
The ABB GoFa [37] is a new line of cobots with three kinematic variants: 5 kg, 10 kg, and 12 kg. The 5 kg and 12 kg variants have intersecting axes $(1,2)$. All variants have parallel axes $(2,3)$, intersecting axes $(4,5)$, and nonconsecutive intersecting axes $(4,6)$. We found examples of non-singular changes of IK solutions for all three variants. To the best of our knowledge, we are the first to report that the GoFa is a cuspidal manipulator.
@@
Although this example exceeds the robot joint limits, we can also find other examples within the joint limits and verify the motion in RobotStudio.
@@ 3.4 Robots with Three Parallel Axes May Be Cuspidal
Others have mentioned that robots with three parallel axes are noncuspidal [7]. Although this is true for common robots such UR cobots which have other intersecting axes and can be solved with a series of quadratics, this is false in general. One can easily find a counterexample by searching through random robots and poses and testing if any MoveJ path is nonsingular. For example, the robot with kinematic parameters
@@ 4 Path Planning Using Graph Method
While the IK problem finds the joint angles for a single end effector pose, the path planning problem extends this to find the entire joint path for an end effector path. Applying IK to each point in the end effector path is not sufficient because the joint path must be continuous and may have other constraints, and the IK problem admits multiple solutions.
@@ 4.1 Problem Formulation
Denote an end effector path $\underline{\chi}_{0T}(\lambda)$ with path index $\lambda\in[0,L]$, and denote a point along the path by $\chi_{0T}(\lambda)$. The goal of the path planning problem is to find the optimal continuous corresponding joint-space path $\underline{q}(\lambda)$. If desired, the path can be indexed to time with a trajectory $\underline{\lambda}(t)$, where $t\in[0,T]$. We can discretize the path into $K+1$ evenly spaced points so that $L=K\Delta\lambda$. Define the discrete path index $k\in\{0,1,\dots,K\}$ so that $\chi_{0T}[k]=\chi_{0T}(k\Delta\lambda)$. Similarly, we can represent the joint-space solutions using the discretized path $\underline{q}[k]$.
@@ 4.2 Graph Method
For a 6-DOF manipulator, the task-space pose $\chi_{0T}[k]=(R_{0T}[k],p_{0T}[k])$ consists of the orientation and position of the end effector task frame in the robot base frame. For each $\chi_{0T}[k]$ the IK solutions are $q^{1}[k],\ q^{2}[k],\ \dots,\ q^{M[k]}[k]$, where $M[k]$ is the number of IK solutions for $\chi_{0T}[k]$. (A special case is when the IK solutions include a solution at an internal singularity, in which case self-motion is possible at that solution.) We construct a weighted directed acyclic graph in the following manner:
@@
  • Each IK solution vertex $V_{k,m}$ has a weighted edge to $V_{k+1,l}$ if the metric $c(q^{l}[k+1],q^{m}[k]))$ is less than some user-defined $\epsilon=\Delta\lambda\epsilon_{0}$. If the optimization metric is (9) and $\dot{q}_{max}$ is the maximum desired joint velocity, then the threshold $\epsilon_{0}$ can be set according to $\epsilon_{0}\approx\left\|{\dot{q}_{max}}\right\|^{2}$.
  • Vertex $S$ is connected to each $V_{0,m}$, and vertex $F$ is connected to each $V_{K,m}$ with edges of weight 0 directed from $S$ and to $F$.
@@
Some care must be taken to properly tune the sampling density $\Delta\lambda$ and the metric threshold $\epsilon_{0}$. There must be enough samples to reasonably approximate the joint-space paths as piecewise linear approximations, and the threshold must be set to correctly discriminate between connected and disconnected IK solutions.
@@
Limits on certain joint angles means that we need to take into account the total number of turns rather than only considering $q_{i}\in[-\pi,\pi]$. Care must be taken for cyclical paths so that repeating the path does not cause the joint to exceed the joint limit. Velocity or acceleration constraints can be used to determine if two vertices are connected with an edge (hard constraint) or they can be incorporated into the cost of each edge (soft constraint).
@@
A joint-space path encounters a singularity if $\det(J)=0$ for some point along the path. If edges are not made between vertices with different $\operatorname{sign}(\det(J))$ then the path planner avoids returning singular paths.
@@
By finding all connections between nodes at $k=0$ and $k=K$ (which represent the same IK solutions), we can make a new directed graph showing connections between IK solutions. Then, we must find graph cycles starting and ending at the same IK solution.
Alternatively, rather than constructing a new graph, we can instead connect each $V_{0,m}$ to the corresponding $V_{K,m}$ and find minimum-weight cycles directly.
@@
To account for missed IK solutions, also connect vertex $V_{k,m}$ to vertex $V_{k+2,l}$ if the metric
@@
is less than $\epsilon$. Similarly, $S$ and $F$ should be connected to $V_{1,m}$ and $V_{K-1,m}$. If desired, nodes more than 2 timesteps away may be connected similarly.
Due to the triangle inequality, paths with fewer edges and with larger $\Delta\lambda$ will have a smaller weight than paths between the same vertices with more edges and with smaller $\Delta\lambda$. To reduce the error in calculating the total path cost, edges should be added to the graph in multiple passes. First, only connect vertices with distance $\Delta\lambda$. Then, connected vertices with distance $2\Delta\lambda$ only if they are not already connected.
Another strategy to prevent missed IK solutions when using IK-Geo is to utilize continuous approximate solutions. This allows for Cartesian paths at or even through boundary singularities. This means paths can be at the edge of the workspace or can switch between aspects. Care should be taken to ensure the task-space path deviation is not excessive when using these solutions. Continuous approximate solutions are also useful as part of IK based on 1D search. In this case, continuous approximate solutions may occur when the error function has a zero at the edge of its branch, which is where the error becomes vertical. These continuous approximate solutions at the edge of the error function branch are usually not singular IK solutions.
@@ 4.3 Evaluation
An example of the graph method for planning a straight-line Cartesian move for the FANUC CRX-10iA/L is shown in Fig. 8. This example uses 10 equally-spaced sample points to make the graph easy to see, but in practice, the density of sample points should be higher. The subtle gray background plot in Fig. 8 was created using a higher density of 100 sample points.
The plots shown in Figs. 2, 3, and 4 were constructed using the graph method. All three examples are for straight-line MoveL paths. There are 100 sample points in Fig. 2(b) and 200 sample points in Figs. 3 and 4.
In Section 5, we will also demonstrate using the graph method inside an optimization loop. Rather than a simple straight-line path, these examples will use more complicated helical paths each with 500 equally-spaced samples.
@@ 5 Path Optimization
Robots commonly perform operations such as welding where there is freedom in how to fixture the workpiece with respect to the robot, either by moving the workpiece or by moving the robot. In such a scenario, a natural problem to solve is how to find the optimal pose of the workpiece based on some optimization metric.
@@ 5.2 Solution Method
A locally optimal workpiece frame pose $\chi_{0P}$ can be found using an optimization method such as fminsearch() in MATLAB [39], which is a derivative-free local optimizer. To find a feasible starting point, randomly search over $\mathrm{SO}(3)$ and $\mathbb{R}^{3}$ based on the robot workspace until the graph-based path planner finds a feasible $\underline{q}[k]$. Then, optimization is performed with the graph-based path planner running at each iteration. If desired, the global minimum can be found with high probability if the optimization is performed with multiple random starting points.
Represent the rotation $R_{0P}$ by a unit quaternion with scalar part $\cos(\theta/2)$ and vector part $k\sin(\theta/2)$, where $k$ is the axis of rotation and $\theta$ is the rotation angle. During optimization, the unit-length constraint does not need to be enforced and the quaternion can be normalized before use. However, it should be verified after optimization that the norm does not become too small or large.
@@
If there are no joint limits or position constraints that depend on the first joint angle $q_{1}$ then rotating the path about the first joint axis $h_{1}$ does not change the optimization.
@@
For the above analysis to be valid, we must show that any rotation $R$ can be represented as a rotation about a vector in the $xy$ plane followed by a rotation about $z$, that is $R=R_{z}R_{xy}$. Decompose the rotation into $z$$x$$z$ Euler angles:
@@ 5.3 Evaluation
Optimization was performed for both a 3R and 6R robot using a helical toolpath consisting of 500 equally-spaced sample points for the path variable $\lambda$. The optimization metric was the $L_{2}$ joint velocity norm approximation (9) without any other constraints.
@@
Optimization was also performed for the FANUC CRX-10iA/L with a 6-DOF helical toolpath and two initial guesses. Results are shown in Figs. 9(b), 10(b), and 11(b). RMS joint velocity over path length decreased from 13.9328 rad/m to 10.0447 rad/m for the optimized pose near $z=0.75$ and decreased from 13.1054 rad/m to 9.9929 rad/m for the optimized pose near $z=-0.5$.
@@ 6 Conclusion
Most robot practitioners are unaware of cuspidal robots, and many of those who are aware suggest cuspidal robots should be avoided. Our suggestion is that cuspidal robots are difficult to use not because of an inherent problem with the kinematics but because there is a current lack of effective IK and path planning algorithms to control them. Indeed, cuspidal robots offer certain advantages. Wrist offsets may lead to simpler mechanical designs and better dynamical performance. Such designs even allow for the upper arm to swing past the lower arm, such as for some models in the CRX line. Having multiple IK solutions available in a single aspect can be beneficial because one IK solution may be preferred over another for obstacle or singularity avoidance, and switching between solutions does not require passing through a singularity. Therefore, it is important not to dismiss cuspidal robots due to their path planning issues but instead to develop new path planning tools so that cuspidal robots can be used to their full potential.
The work presented in this paper is one of the few first steps in improving the path planning capabilities of cuspidal robots. We have shown how to identify cuspidal robots, how to perform path planning given a task-space path, and how to perform path optimization. For the first time, we have shown the ABB GoFa and certain robots with three parallel joint axes are cuspidal. In practice, the identification algorithm can be used to identify the cuspidality of more existing robots and to aid in the design of new robots. The algorithms presented underscore the importance of IK solvers which can efficiently find all IK solutions, such as IK-Geo. In contrast, we showed RobotStudio does not find all solutions for the ABB GoFa. We demonstrated the path planning and optimization algorithms on both straight-line paths as well as more complicated helix paths. The path planning and optimization algorithms have many practical applications for tasks where the entire toolpath is known and the workpiece or robot base can be moved. This includes subtractive processes such as grinding, sanding, and polishing, and additive processes such as welding, spraying, and 3D printing.
@@
  • What is a good metric for the “degree of cuspidality" for a robot? Different robots have different proportions of random poses with non-singular MoveJ changes of IK solution (I.e., the algorithm in Section 3.2 may have a different expected iteration count) and different proportions of random MoveL paths that are infeasible. Such a metric may even be useful for noncuspidal robots. We hope such a metric can help influence the design or selection of robots to prevent cases of infeasible or nonrepeatible paths.
@@
  • Can we apply cuspidality analysis to 7-DOF manipulators? Preliminary testing shows that some redundant serial robot arms parameterized by either one joint angle or by the Shoulder-Elbow-Wrist (SEW) angle are cuspidal.
This work in identification, path planning, and path optimization addresses critical theoretical challenges of cuspidal robots while improving their practical viability. Advancing these methods and addressing the open questions we raise will further unlock the full potential of cuspidal manipulators.